If the average current passing through the secondary coil of the transformer with 40 turns and a secondary voltage of 24-V AC is 10 A, what average current is passing through the primary coil that makes 200 turns and has a primary voltage of 120-V AC?
Answer #1

given
Ns = 40 turns
Vs = 24 V
Is = 10 A
Np = 200 turns
Vp = 120 V
Ip = ?

use, Power supplied to primary coil = power output from secondary coil

Ip*Vp = Is*Vs

Ip = Is*Vs/Vp

= 10*24/120

= 2 A

Note: P stands for primary and S stands fro secondary.